DNA (deoxyribonucleic acid) is the basis of most genes. DNA is made of nucleotides. Each nucleotide can contain any one of these four nitrogenous bases is a linear sequence: A (adenine), G (guanine), C (cytosine), T (thymine). If an experimenter wishes to create a gene with a sequence of 9 bases, where bases can be repeated and order matters, how many different sequences are possible?
Answer by jim_thompson5910(35256)

There are 4 possibilities for the first position

There are 4 possibilities for the second position

...

etc


Now multiply all these out


So there are 4*4*4*4*4*4*4*4*4 = 4^9 = 262144 different sequences